a piece of cloth was carried by wind and seen rising up in the sky and after some time it was suspended in air. state the mathematical expression of the forces acting on the object when suspended in air.
1
Expert's answer
2020-12-14T07:26:55-0500
"F=W-F_{res}\\F=mg-kv^2"